# Understanding Radiation Flow in a Stochastic Medium

### v1

@@ 4.1 Simulation Tools and Initial Conditions
For our calculations, we use conditions close to that of the XFOL and XFLOWS experimental campaigns led by LANL (Johns et al., 2023; Byvank et al., 2024) and the Cassio code developed under LANL’s Advanced Simulation and Computing program. The Cassio code models hydrodynamics using a cell-based adaptive mesh refinement scheme with a two-shock approximate Riemann solver. The code has been verified against a variety of analytic test problems, the most relevant for this problem being the Sedov blast wave (Gittings et al., 2008). Radiation transport can be modeled using a variety of transport solutions including flux-limited diffusion (Gittings et al., 2008), discrete ordinate ($S_{N}$(Fryer et al., 2020a), and Implicit Monte Carlo transport (Urbatsch & Evans, 2006).
@@
The Cassio code has been used extensively in the laboratory experimental community for a wide variety of problems studying hydrodynamics or radiation-hydrodynamics (e.g. Johns et al., 2016; Falk et al., 2017, 2018, 2020a; Fryer et al., 2020b; Falk et al., 2020b; Johns et al., 2021; Coffing et al., 2022; Johns et al., 2023; Fryer et al., 2023). In addition to many of these validation tests, the Cassio has been verified against other experimental codes (e.g. Fatenejad et al., 2013; Falk et al., 2014). It has been applied to a number of radiation-hydrodynamics problems in astrophysics (for a review, see Fryer et al., 2020a) and has been verified against other astrophysical codes as well (Joggerst et al., 2014).
For most of our calculations, we use Cassio’s $S_{N}$ transport scheme with an $N=8$ quadrature set and 71 energy groups. The coarse grid resolution is $0.6\,{\rm\mu m}$ and we use only 2 levels of refinement. We discuss comparisons to other transport schemes and grid resolutions below.
@@
The initial conditions for our simulations are guided by active XFOL and XFLOWS experimental campaigns (Johns et al., 2023; Byvank et al., 2024) that builds upon past Pleiades and COAX radiation flow studies (Fryer et al., 2020b; Johns et al., 2021; Coffing et al., 2022). In these experiments, a laser-driven hohlraum (either at the Omega laser facility in Rochester (Boehly et al., 1995) or the National Ignition Facility at Lawrence Livermore National Laboratory (Haynam et al., 2007) produces a radiation front that then propagates through a target. This experiment consists of a target placed on top of a hohlraum (Figure 2). Lasers fired into the hohlraum create a hot, radiation-dominated plasma that then propagates through the experimental target. The power in this drive peaks quickly and then decays slowly over the course of our few-ns simulation. In our multi-clump, high-drive models, we also consider a simplified flat power (or temperature) for this drive. The target is a silicon dioxide foam filled with vanadium-oxide inclusions; the opacities are obtained from the OPLIB database (Colgan et al., 2016), calculated with the Los Alamos suite of atomic physics codes (Fontes et al., 2015). In this paper, we focus solely on modeling the target, using a radiation source term at its base to represent the hohlraum-produced radiation flow. The target is modeled in cylindrical symmetry with a radius of 0.04 cm and a height of 0.08 cm.
@@
Figure 2: Simplified diagram of the experiment XFOL experiment (Johns et al., 2023) designed to measure radiation flow through an inhomogeneous medium. Lasers heat the gold-line hohlraum cavity, producing a hot plasma that drives a radiation front through a target. Dense inclusions are added into a low-density foam target that is roughly 0.04 cm in size. The radiation front can be probed both using a radiograph to measure the position of the front or a spectrograph to constrain the temperature and composition. The escape of the radiation front from the top of the target can be used to measure the radiation flux versus time.
@@ 4.2 Single Clump Studies: Probing the Fundamental Physics
As a first study, we focus on the effects of clump density. Figure 3 shows the density map of 4 separate simulations 1.6 ns after the launch of the radiation front where the only difference in the 4 simulations is the density of the clump. For these clumps, the composition of clump and flow regions are identical. Only the clump density is varied including models where the clump density is lower and higher than the ambient flow region density. For the low-density clump, radiation rushes through the cavity, leading the flow in the ambient region. For high density clumps, the radiation flows around the clumps. For modest density increases (e.g. 10 times the ambient medium), the radiation can both ablate and compress the clump. However, the timescale to do this increases with increasing clump density and, at the snapshot in time shown in Figure 3, the radiation flow has little effect on the clump.
Figure 3: Density map for 4 cylindrical models after 1.6 ns where the clump density is varied (left to right) from 1/2 of the flow region density to 100 times the flow region density. The opacity varies dramatically as the radiation front sweeps down the flow region. For the cold material, the initial Rosseland opacity for the low-energy X-ray photons in this flow is $>4\times 10^{4}{\rm\,cm^{2}\,g^{-1}}$. Once heated, the opacity drops by over an order of magnitude and the mean free path in the foam region is roughly 0.01 cm and the flow region is roughly 8 mean free paths long. The clump opacity is roughly 50% higher. In the high density clump (roughly the density used for most of our calculations), the mean free path is 0.03–0.1 $\mu$m and since the bulk of the clump remains cold throughout the calculation, this mean free path remains low.
The radiation heats up the clump, striving to place it in temperature equilibrium. If the clump is denser than that of its surroundings, it begins to blow a wind, sending a material flow back into the flow region. As we shall discuss below, if the composition of the clumps is different than the ambient medium, these outflows mix into the flow region. If this material has higher opacity, it can choke off the radiation flow. Another features of these outflows is that they can cause shocks, raising the temperature of the outflow (Figure 4). These shocks are critical in understanding supernova outflows, producing extended Ultraviolet and X-ray emission (Fryer et al., 2020a). The alterations in the flow caused by a low-density region can also alter the temperature profile. Especially in ambient media where the radiation is in the transport or free-streaming regime, this heating from radiation-matter interactions can dramatically alter the emerging spectra. If the radiation flow is highly photon-energy dependent, this effect can also alter the flow of radiation.
Figure 4: Temperature map for 4 cylindrical models after 1.6 ns where the clump density is varied (left to right) from 1/2 of the flow region density to 100 times the flow region density (same as Figure 3).
@@
$P_{\rm clump/flow}=\rho_{\rm clump/flow}/\langle A\rangle RT_{\rm clump/flow}+a_{\rm rad}/3T^{4}_{\rm clump/flow}$ (4)
where $P_{\rm clump/flow}$, $\rho_{\rm clump/flow}$ and $T_{\rm clump/flow}$ are the pressure, density and temperature of the clump or flow region, $\langle A\rangle$ is the average atomic weight of the material, $R=8.31\times 10^{7}{\rm erg\,K^{-1}mol^{-1}}$ is the universal gas constant, and $a_{\rm rad}=7.567\times 10^{-15}{\rm\,erg\,cm^{-3}\,K^{-4}}$ is the radiation constant. Figure 5 shows the clump temperature versus the flow-region temperature when the pressure in the clump equals that of the flow region. The lines in this plot denote pressure equilibrium at the two temperatures on the x and y axes. We have set up rough conditions for our experiment where $\rho_{\rm clump/flow}/\langle A\rangle=0.01$ and we vary the clump value up to 30 times higher. At these densities, ideal gas pressure dominates the radiation pressure at the temperatures in the experiment. As the clump is heated, its pressure will exceed that of the flow region and it will expand, blowing a wind into the flow region.
Figure 5: Curves of pressure equilibrium versus the temperature of the flow region (x axis) and the temperature in the clump (y axis). In this scenario, we assume $\rho_{\rm flow}/\langle A\rangle$ for the flow region is $0.01{\rm\,mol\,cm^{-3}}$ and vary the quantity $n=\rho_{\rm clump}/\langle A\rangle$ over four different values. These conditions are reasonably close to those in our laboratory experiment.
@@
$T_{\rm crit}=a_{\rm rad}/3T^{4}_{\rm flow}/(\rho_{\rm clump}/\langle A\rangle R)\,.$ (5)
Table 1 provides rough values of the critical temperature above which radiation pressure dominates and this blow-off is minimized. For the different phenomena, both the clump density and temperature of the flow region can vary by many orders of magnitude. For most stars, high stellar radii and low mass-loss rates lead to conditions where the stellar radiation dominates the pressure. Even though Wolf-Rayet stars are hotter, the mass loss rates are higher and the stars are more compact. Near the stellar surface, material pressure can be important.
Phenomena $\rho/\langle A\rangle$ $T_{\rm crit}$ $T_{\rm flow}$ Dominant Pressure
@@
From our analytic models, we expect the outflows to depend upon both the specific heat and opacity of the material. Figure 6 shows the temperature map from two simulations at two snapshots in time where the specific heat is raised and lowered by an order of magnitude. For high specific heat models, the temperature in the clump need only rise modestly to achieve sufficient pressures to push out against the ambient medium, driving an outflow into the flow region. If the clump material has a lower specific heat, it must become hotter to achieve the requisite pressures to drive an outflow. The outflow will take longer to expand and will be hotter.
Figure 6: Temperature map for two cylindrical models after 1.7 and 2.0 ns where the specific heat of the clump is lowered and raised by an order of magnitude.
Figure 7 shows the density and volume fraction ($f_{\rm vol}$) maps corresponding to the temperature maps in Figure 6. Because the outflow starts earlier for the higher specific heat model, it expands more quickly into the ambient medium. In addition to being lower temperature, the density of the outflow is lower for higher $C_{v}$ clump. If the opacity of this material is high, it can cut off the radiation flow. Although the high specific clump material expands more quickly, it is less dense, so the impact on the radiation flow may be greater for the clumps with lower specific heats.
@@
Figure 8: Density and temperature maps for two cylindrical models at 2.0 ns where the opacity of the material in the clump is lowered and raised by an order of magnitude.
Figure 9: Opacity and volume fraction maps for two cylindrical models at 2.0 ns where the opacity of the material in the clump is lowered and raised by an order of magnitude.
One important aspect of many computational studies of transport in inhomogeneous media is that all opacity databases that we are aware of assume that the material is a plasma. Especially for laboratory experiments, the embedded clumps are in a solid phase, and the opacity of this material can be very different from opacities obtained under plasma conditions. Electrons obey Fermi-Dirac statistics, which ensures that at solid densities, the partial occupation of higher-lying energy levels decays slowly, resulting in the long characteristic Fermi-Dirac tail. However in dense plasmas, the eigenspectrum is continuous and the partial occupations die out rapidly. This behavior governs the intra (inter) band transitions, or equivalently, bound–bound, bound–free, and free–free electron transitions, and therefore has a direct impact on optical spectra Filinov et al. (2003); Helled et al. (2020). A recent study of asymmetric carbon–hydrogen (CH) mixtures has shown that compressing the system to three times its density at $k_{B}T=10$ eV results in an almost threefold increase in its electrical and thermal conductivities White et al. (2023). A compressed “nonideal” plasma state emerges in numerous dynamic compression environments such as in the laser irradiation of condensed matter in pinched electric discharge experiments, or when the critical density values are attained at high static pressures Boehler (1993); Aquilanti et al. (2015).
@@
In these studies, we have assumed numerical issues (e.g. grid resolution, radiation transport methods) do not significantly alter our results. Our standard simulations are fairly high resolution (in a $0.04\,{\rm cm}=400\,\mu m$ target, our resolution is $0.6\,\mu m$). However, ideally, our calculations would resolve the mean free path of the radiation and typical mean free paths in our clump are $0.15\,\mu m$. To test the resolution in our standard calculations, we ran a series of high-resolution calculations where the resolution was $0.15\,\mu m$. The results, comparing the standard- to high-resolution calculations is shown in Figure 10. Although small structures seem to be appearing at $t=2\,$ns, the outflow and shock positions are identical in these calculations.
Figure 10: Density and temperature maps for two cylindrical models at 2.0 ns where the resolution is varied by a factor of four (16 times more zones in 2-dimensions).
@@
Figure 11: Density and temperature maps for our standard cylindrical model at 2.0 ns using Implicit Monte Carlo. Compare these results to Figure 10. Despite mild differences in initial conditions and large differences in the numerical approaches, the interaction with the clump with these two schemes is very similar.

### v3

@@ 3.2 Experiments and Scaling
These experiments have demonstrated the potential of laboratory experiments to probe the different physical aspects of the inhomogeneous radiation flow problem. As increasingly sophisticated diagnostics are developed, these experiments will provide a means to test radiation flow and its hydrodynamic feedback.
@@ 4.1 Simulation Tools and Initial Conditions
For our calculations, we use conditions close to that of the XFOL and XFLOWS experimental campaigns led by LANL (Johns et al., 2023; Byvank et al., 2024) and the Cassio code developed under LANL’s Advanced Simulation and Computing program  (for details, see Gittings et al., 2008; Urbatsch & Evans, 2006; Fryer et al., 2020a). The Cassio code models hydrodynamics using a cell-based adaptive mesh refinement scheme with a two-shock approximate Riemann solver using the methods described in the xRAGE code (Gittings et al., 2008). The code has been verified against a variety of analytic test problems, the most relevant for this problem being the Sedov blast wave (Gittings et al., 2008). Cassio has several implementations of multi-material physics. For our calculations, we use the default prescription. This prescription assumes that when multiple materials are in a zone, the materials are in pressure-temperature equilibrium. The relative density of the materials is determined by the specific heat capacities of the materials. In this prescription, even if the mass fractions are the same, the materials can have different volume fractions and different densities. The inferred densities are used in calculating the opacities. Multi-material solutions such as this are not commonly used in astrophysics (for monatomic gases, the specific heat capacities are all 3/2), but are critical in a lot of applied physics, experiments and engineering solutions (Huhn et al., 2025). Time-dependent radiation transport can be modeled using a variety of transport solutions including flux-limited diffusion (Gittings et al., 2008), discrete ordinate ($S_{N}$(Fryer et al., 2020a), and Implicit Monte Carlo transport (Urbatsch & Evans, 2006) using a multi-group energy resolution.
@@
The Cassio code has been used extensively in the laboratory experimental community for a wide variety of problems studying hydrodynamics or radiation-hydrodynamics (e.g. Johns et al., 2016; Falk et al., 2017, 2018, 2020a; Fryer et al., 2020b; Falk et al., 2020b; Johns et al., 2021; Coffing et al., 2022; Johns et al., 2023; Fryer et al., 2023). In addition to many of these validation tests, Cassio has been verified against other experimental codes (e.g. Fatenejad et al., 2013; Falk et al., 2014). Cassio and its related flux-limited diffusion version xRAGE have been applied to a number of radiation-hydrodynamics problems in astrophysics, mostly studying supernova light curves (Fryer et al., 2009, 2010a, 2010b; Whalen et al., 2013a, b, e; Frey et al., 2013a, b; Whalen et al., 2013d, c, f, g, 2014a, 2014b; Smidt et al., 2014, 2015; Fryer et al., 2020a) and has been verified against other astrophysical codes as well (Joggerst et al., 2014).
Our calculations used a binned opacity approach similar to that used in many supernova codes including SuperNu (Wollaeger et al., 2013) and STELLA (Bartunov & Blinnikov, 1992). The binned approach has been compared to other line sampling and expansion opacity methods (Wollaeger & van Rossum, 2014; Fontes et al., 2020; Blondin et al., 2022) with reasonable agreement. The binned opacities are obtained from the OPLIB database (Colgan et al., 2016) calculated with the Los Alamos suite of atomic physics codes (Fontes et al., 2015). The bins are created using a Rosseland mean approach, but a Planck averaging can be used (and is sometimes implemented) in the emission terms of the transport equation. For most of our calculations, we use Cassio’s discrete-ordinate, $S_{N}$, transport scheme with an $N=8$ quadrature set and 71 energy groups [^2] For more details on the $S_{N}$ method implemented in Cassio, see Carlson et al. (1965); Fryer et al. (2020a). The transport solution is placed on the same grid as the hydrodynamics calculations. The coarse grid resolution is $0.6\,{\rm\mu m}$ and we use only 2 levels of refinement. We discuss comparisons to other transport schemes and grid resolutions below.
@@
The initial conditions for our simulations are guided by active XFOL and XFLOWS experimental campaigns (Johns et al., 2023; Byvank et al., 2024) that builds upon past Pleiades and COAX radiation flow studies (Fryer et al., 2020b; Johns et al., 2021; Coffing et al., 2022). In these experiments, a laser-driven hohlraum (either at the Omega laser facility in Rochester (Boehly et al., 1995) or the National Ignition Facility at Lawrence Livermore National Laboratory (Haynam et al., 2007) produces a radiation front that then propagates through a target. This experiment consists of a target placed on top of a hohlraum (Figure 2). Lasers fired into the hohlraum create a hot, radiation-dominated plasma that then propagates through the experimental target. The power in this drive peaks quickly and then decays slowly over the course of our few-ns simulation. In our multi-clump, high-drive models, we also consider a simplified flat power (or temperature) for this drive. The target is a silicon dioxide foam filled with vanadium-oxide inclusions. In this paper, we focus solely on modeling the target, using a radiation source term at its base to represent the hohlraum-produced radiation flow. The radiation source is a fit to a particular set of hohlraum models with a radiation temperature that is relatively flat over the course of the simulation (peaking at 150eV at 1 ns). The target is modeled in cylindrical symmetry with a radius of 0.04 cm and a height of 0.08 cm.
@@
Figure 2: Simplified diagram of the experiment XFOL experiment (Johns et al., 2023) designed to measure radiation flow through an inhomogeneous medium. Lasers heat the gold-line hohlraum cavity, producing a hot plasma that drives a radiation front through a target. Dense inclusions are added into a low-density foam target that is roughly 0.04 cm in size. The radiation front can be probed both using a radiograph to measure the position of the front or a spectrograph to constrain the temperature and composition. The escape of the radiation front from the top of the target can be used to measure the radiation flux versus time (the current version of this experiment does not have a detector for this measurement).
@@ 4.2 Single Clump Studies: Probing the Fundamental Physics
As a first study, we focus on the effects of clump density. Figure 3 shows the density map of 4 separate simulations 1.6 ns after the launch of the radiation front where the only difference in the 4 simulations is the density of the clump. For these clumps, the composition of clump and flow regions are identical. Only the clump density is varied including models where the clump density is lower and higher than the ambient flow region density. For the low-density clump, radiation rushes through the cavity, leading the flow in the ambient region. For high density clumps, the radiation flows around the clumps. For modest density increases (e.g. 10 times the ambient medium), the radiation can both ablate and compress the clump. However, the timescale to do this increases with increasing clump density and, at the snapshot in time shown in Figure 3, the radiation flow has little effect on the high-density clump.
[…]
The radiation heats up the clump, striving to place it in temperature equilibrium. If the clump is denser than that of its surroundings, it begins to blow a wind, sending a material flow back into the flow region. As we shall discuss below, if the composition of the clumps is different than the ambient medium, these outflows mix into the flow region. If this material has higher opacity, it can choke off the radiation flow. Another feature of these outflows is that they can cause shocks, raising the temperature of the outflow (Figure 4). These shocks are critical in understanding supernova outflows, producing extended Ultraviolet and X-ray emission (Fryer et al., 2020a). The alterations in the flow caused by a low-density region can also alter the temperature profile. Especially in ambient media where the radiation is in the transport or free-streaming regime, this heating from radiation-matter interactions can dramatically alter the emerging spectra. If the radiation flow is highly photon-energy dependent, this effect can also alter the flow of radiation.
Figure 4: Effective radiation temperature map for 4 cylindrical models after 1.6 ns where the clump density is varied (left to right) from 1/2 of the flow region density to 100 times the flow region density (same as Figure 3). Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
@@
$P_{\rm clump/flow}=n_{clump/flow}RT_{\rm clump/flow}+\frac{a_{\rm rad}}{3}T^{4}_{\rm clump/flow}$ (4)
where $P_{\rm clump/flow}$, $T_{\rm clump/flow}$ and $n_{\rm clump/flow}$ are the pressure, temperature and number density of the clump or flow region, where the number density is given by the density ($\rho_{\rm clump/flow}$) divided by the average atomic weight of the material ($\langle A\rangle$), $R=8.31\times 10^{7}{\rm erg\,K^{-1}mol^{-1}}$ is the universal gas constant, and $a_{\rm rad}=7.567\times 10^{-15}{\rm\,erg\,cm^{-3}\,K^{-4}}$ is the radiation constant. Figure 5 shows the clump temperature versus the flow-region temperature when the pressure in the clump equals that of the flow region. The lines in this plot denote pressure equilibrium at the two temperatures on the x and y axes. We have set up rough conditions for our experiment where the number density, $n_{clump/flow}$, =0.01=0.01 and we vary the clump value up to 30 times higher. At these densities, ideal gas pressure dominates the radiation pressure at the temperatures in the experiment. As the clump is heated, its pressure will exceed that of the flow region and it will expand, blowing a wind into the flow region.
Figure 5: Curves of pressure equilibrium versus the temperature of the flow region (x axis) and the temperature in the clump (y axis). In this scenario, we assume the number density, $n_{flow}$, for the flow region is $0.01{\rm\,mol\,cm^{-3}}$ and vary the number density of the clump, $n_{clump}$, over four different values. These conditions are reasonably close to those in our laboratory experiment.
@@
$T_{\rm crit}=\frac{a_{\rm rad}T^{4}_{\rm flow}}{3n_{clump}R}\,.$ (5)
Table 1 provides rough values of the critical temperature above which radiation pressure dominates and this blow-off is minimized. For the different phenomena, both the clump density and temperature of the flow region can vary by many orders of magnitude (Debnath et al., 2024). For most stars, high stellar radii and low mass-loss rates lead to conditions where the stellar radiation dominates the pressure. Even though Wolf-Rayet stars are hotter, the mass loss rates are higher and the stars are more compact (Crowther, 2007; Sander & Vink, 2020). Many stellar models predict bursts of mass ejection from the star, either through explosive shell burning, opacity-driven instabilities and pressure waves (Fryer et al., 2006; Herwig et al., 2014; Quataert et al., 2016). Especially in stripped stars, these can produce prompt ejection rates that are much higher than the mass loss rates of WR stars. We include all of these scenarios in Table 1.
Phenomena $n=\rho/\langle A\rangle$ $T_{\rm crit}$ $T_{\rm flow}$ Dominant Pressure
@@
From our analytic models, we expect the outflows to depend upon both the specific heat and opacity of the material. Especially if the the denser clumps are caused through common envelope mass ejection, the clumps are likely to have a different composition than the surrounding wind material. Although the variation in the specific heat is less important in astrophysical plasmas than in laboratory experiments, opacity differences can be an order of magnitude. Figure 6 shows the temperature map from two simulations at two snapshots in time where the specific heat is raised and lowered by an order of magnitude. For high specific heat models, the temperature in the clump need only rise modestly to achieve sufficient pressures to push out against the ambient medium, driving an outflow into the flow region. If the clump material has a lower specific heat, it must become hotter to achieve the requisite pressures to drive an outflow. The outflow will take longer to expand and will be hotter.
Figure 6: Effective radiation temperature map for two cylindrical models after 1.7 and 2.0 ns where the specific heat of the clump is lowered and raised by an order of magnitude. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
Figure 7 shows the density and volume fraction ($f_{\rm vol}$) maps [^3] The volume fraction is close to the mass fraction in each zone. However, Cassio assumes different materials in a single zone are in Pressure and Temperature equilibrium and the different specific heats of the different materials will lead to different material densities and volume fractions that are slightly different than the mass fractions in a given zone. corresponding to the temperature maps in Figure 6. Because the outflow starts earlier for the higher specific heat model, it expands more quickly into the ambient medium. In addition to being lower temperature, the density of the outflow is lower for higher $C_{v}$ clump. If the opacity of this material is high, it can cut off the radiation flow. Although the high specific heat material expands more quickly, it is less dense, so the impact on the radiation flow may be greater for the clumps with lower specific heats.
@@
Figure 8: Density and effective radiation temperature maps for two cylindrical models at 2.0 ns where the opacity of the material in the clump is lowered and raised by an order of magnitude. As the opacity of the clump increases, the radiation is unable to penetrate the clump leading to a lower-mass wind. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
Figure 9: Opacity and volume fraction maps for two cylindrical models at 2.0 ns where the opacity of the material in the clump is lowered and raised by an order of magnitude. For the higher-opacity simulation, the mass of the wind is lower, but the opacity of this wind can be much higher. This will alter the radiation flow.
One important aspect of many computational studies of transport in inhomogeneous media is that all opacity databases that we are aware of assume that the material is a plasma. Especially for laboratory experiments, the embedded clumps are in a solid phase, and the opacity of this material can be very different from opacities obtained under plasma conditions. Electrons obey Fermi-Dirac statistics, which ensures that at solid densities, the partial occupation of higher-lying energy levels decays slowly, resulting in the long characteristic Fermi-Dirac tail. However in dense plasmas, the eigenspectrum is continuous and the partial occupations die out rapidly. This behavior governs the intra (inter) band transitions, or equivalently, bound–bound, bound–free, and free–free electron transitions, and therefore has a direct impact on optical spectra (Filinov et al., 2003; Helled et al., 2020). A recent study of asymmetric carbon–hydrogen (CH) mixtures has shown that compressing the system to three times its density at $k_{B}T=10$ eV results in an almost threefold increase in its electrical and thermal conductivities (White et al., 2023). A compressed “nonideal” plasma state emerges in numerous dynamic compression environments such as in the laser irradiation of condensed matter in pinched electric discharge experiments, or when the critical density values are attained at high static pressures (Boehler, 1993; Aquilanti et al., 2015). When using laboratory experimental data to test our hydrodynamic feedback to radiative heating, we must understand this physics.
@@
In these studies, we have assumed numerical issues (e.g. grid resolution, radiation transport methods) do not significantly alter our results. Our standard simulations are fairly high resolution (in a $0.04\,{\rm cm}=400\,\mu m$ target, our resolution is $0.6\,\mu m$). However, ideally, our calculations would resolve the mean free path of the radiation. Typical mean free paths in our clump are $0.15\,\mu m$. To test the resolution in our standard calculations, we ran a series of high-resolution calculations where the resolution was $0.15\,\mu m$. The results, comparing the standard- to high-resolution calculations is shown in Figure 10. Although small structures seem to be appearing at t=2t=2\,ns, the outflow and shock positions are identical in these calculations.
Figure 10: Density and effective temperature maps for two cylindrical models at 2.0 ns where the resolution is varied by a factor of four (16 times more zones in 2-dimensions). Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
@@
Figure 11: Density and effective radiation temperature maps for our standard cylindrical model at 2.0 ns using Implicit Monte Carlo. Compare these results to Figure 10. Despite mild differences in initial conditions and large differences in the numerical approaches, the interaction with the clump with these two schemes is very similar. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
@@
Our single-clump observations demonstrate the importance of hydrodynamic feedback in studies of inhomogeneous radiation flow. In many cases, the clump material can have a different composition than the flow region. If the clump opacity is very different than the flow region, the blow-off from the heating clump can drastically alter the opacity. The different line features in the different materials make it possible for laboratory experiments to test the extent of the blow-off. In supernovae, this blow-off can explain the sudden appearance of lines in the supernova spectra.